OpenStudy (anonymous):

Given the slope of a line is 2 and it goes through the point `(2, 1)`, what is the y-intercept? A. 2 B. 1 C. - 2 D. - 3

OpenStudy (tkhunny):

Have you considered substituting? y = mx+b y = 2x+b 1 = 2(2)+b Now what?